In a right triangle , find the length of the side not given.
b=1,c=sqrt26
The length of the third side is

is it 5?

Answer by mananth(16949)

sqrt26^2-1^2= leg^2
26-1= leg^2
25=leg^2
5=leg
...
you are right
..
m.ananth@hotmail.ca